OpenStudy (anonymous):

write an equation of the line that passes through (1,-2) and is perpendicular to the line that passes through (4,2) and (0,4).

OpenStudy (anonymous):

First, find the slope of the line passing through (4,2) and (0,4). Slope = difference in y-values/difference in x-values slope = (4-2)/(0-4) = -1/2 The line perpendicular to this line must have a slope that is the negative reciprocal of - 1/2. So the negative reciprocal of -1/2 is 2. So all what you need to do, is to find the equation of the line with slop 2 and passing through (1,-2). I am sure you can do that.